OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of

OpenStudy (radar):

please complete the sentence. parallel to a line with slope of........

OpenStudy (anonymous):

sorry it's 3/4

OpenStudy (radar):

O.K then we know that the line for which we are to develop the standard form equation is parallel so it too will have an equal slope of 3/4

OpenStudy (radar):

You can use the equation for slope. m=(y-y1)/(x-x1). plug in 3/4 for m. 3/4=(y-(-1))/(x-2) that works out to be: 3x-6=4y+4 placing in standard form: 3x-4y-10=0

OpenStudy (anonymous):

do you think you could also help me with this one? Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (radar):

Yes, you can use the same formula: m=(y1-y2)/(x1-x2) this time m is unknown but the two points are known. Thus m=(6-5)/(4-1)=1/3 Now use the formula with the new found value of m and use just one point I would use G. 1/3=(y-6)/(x-4) this becomes: x-4=3y-18 placing in standard form x-3y+14=0